Please Help: How much do you need to set aside in an account that pays 4.9% interest compounded semi-annually to have 75,000.00 in 15 years?
Answer by scott8148(6628)

semi-annually means two compoundings per year; each at half the yearly interest rate

a=p(1+r)^n

a is 75000 __ p is the principle (what we're trying to find) __ r is 4.9%/2 or .0245 __ n is 15*2 or 30

75000=p(1.0245)^30 __ dividing by (1.0245)^30 __ 7500/[(1.0245)^30]=p __ 36282.93=p